Numbers 1,2,3,4,5,6,7,8,9 are written on seperate cards. The cards are shuffled and the top one turned over. Let A = to an even number. B = to a number greater than 6. Which one of the following statements is incorrect?
1. The sample space is S = {1,2,3,4,5,6,7,8,9}
2. P(A) = 4/9
3. P(B) = 1/3
4. P(A and B) = 1/9
5. P(A or B) = 7/9

Answer by jrfrunner(365)

1 thru 4 are correct
#5 is wrong P(A or B) =P(A) + P(B)-P(A and B) = 4/9+1/3-1/9 = 6/9 or 2/3